# Neural Uncertainty Principle: A Unified View of Adversarial Fragility and LLM Hallucination

### 20 Mar 2026 v1

@@ Results
Table III shows that Risk-Cos achieves a high Top-1 Hit Rate (0.76) while also attaining the lowest Mean Regret (0.188) and the strongest Best-vs-Rest AUROC (0.604) among all methods. Notably, Risk-Margin matches the same Top-1 Hit Rate (0.76) but exhibits substantially higher regret (0.325) and near-chance Best-vs-Rest AUROC (0.481). This discrepancy indicates that Risk-Margin can occasionally identify the single best prompt, but its overall ordering of the five variants is unstable: when it fails to hit the best, it tends to select a much worse variant, increasing regret and collapsing pooled discrimination toward chance.
Robustness to judge-consensus thresholds
Table IV reports Best-vs-Rest AUROC under three consensus buckets. Risk-Cos remains consistently above chance across all buckets (Hard/Medium/All), whereas Risk-Entropy and Risk-Loss (NLL) remain weak or unstable. This pattern aligns with Exp. V-F: under the prefill-only constraint, prompt entropy and prompt NLL mainly reflect how predictable the prompt wording is under the model and do not reliably predict downstream solution quality.
Overall, Exp. V-G shows that Risk-Cos provides a practical, decoding-free criterion for selecting among semantically equivalent prompt formulations. Compared with other prefill-only baselines, Risk-Cos yields similar top-1 hit rate but a more stable ranking, resulting in lower regret.
@@ VI-B Unified interpretation across vision and language
Within this limit-based view, adversarial fragility and hallucination become two opposite ways of mismanaging the same uncertainty budget. Table V summarizes how NUP manifests differently across modalities.
TABLE V: Two regimes from the NUP. The NUP constrains an effective conjugate volume $V=\Delta x\,\Delta p\,\sqrt{1-\rho^{2}}\geq\kappa$, where $\rho$ is the operator correlation between $\hat{x}_{u}$ and $\hat{p}_{u}$ in the RS inequality. Empirically, we diagnose the input–gradient coupling channel entering $\rho$ using the $C$-Probe: in vision we use $|\cos(x,p)|$ (inputs are normalized), while in LLM prompting we use the centered $|\cos(\bar{x},\bar{p})|$.
@@ VI-D Limitations and future directions
Our analysis is most informative when there exists a nontrivial boundary-relevant subset that the loss-phase construction can emphasize. In near “pure fitting” regimes—e.g., when the loss is essentially zero on the relevant support, or the input–output mapping behaves close to bijective with little effective compression [40]—the weighting $w_{c}(x)\propto\mathcal{L}_{c}(x)^{2}$ provides little contrast and distinct loss-induced states $\psi_{c}$ are not meaningfully separated; consequently, the notion of a stressed boundary layer becomes ill-posed and the CC-Probe may lose discriminative power. In addition, the probe is loss-dependent (CE in vision; shifted prompt NLL in language), and Exp. 3 indicates robustness can differ across attack losses (CE vs. DLR), motivating multi-loss evaluation (e.g., APGD-CE and APGD-DLR) and defenses that cover multiple geometry channels. Finally, CC-Probe is a proxy rather than a direct estimator of $\rho_{c}(u)$ or slack $S_{c}(u)$, and prefill-only LLM use still requires one backward pass on prompt embeddings. Future work includes designing alternative state constructions when $\mathcal{L}_{c}$ is near-degenerate, developing stable estimators of direction-dependent slack, and extending coupling-based monitoring/control to multimodal and preference-optimized settings.
@@ VII Conclusion
We have introduced the Neural Uncertainty Principle (NUP), a fundamental geometric constraint showing that input projections and their loss gradients are conjugate observables under a loss-induced state. NUP unifies adversarial vulnerability in vision and hallucination in LLMs as opposite extremes on the same conjugate plane—saturation versus slack. This principle yields a practical single-backward probe (CC-Probe) that diagnoses boundary anomalies across modalities and guides targeted interventions: ConjMask/LogitReg for robust vision, and prefill-stage risk scoring for hallucination detection and prompt selection. By reframing two isolated reliability challenges as a single conjugate trade-off, NUP provides insights to the pursuit of robust and faithful AI.

### 26 Mar 2026 v2

@@ Results
Table III shows that Risk-Cos achieves the highest Top-1 Hit Rate (0.83) and the lowest Mean Regret (0.433) among all compared prefill-only risk scores. This indicates that, under the prefill-only constraint, Risk-Cos is the strongest criterion on the two primary selection metrics reported in the main text.
Risk-Margin attains a similarly high Top-1 Hit Rate (0.81), but its Mean Regret is larger (0.591). This gap suggests that although Margin often selects a competitive prompt, its failures are more costly when it misses the judge-best variant. By comparison, Risk-Loss (NLL) and Risk-Entropy perform worse on these two primary metrics, with lower hit rates (0.78 and 0.77, respectively) and higher regret (0.615 and 0.636, respectively).
Judge reliability
The judge panel exhibits strong agreement on the binary hallucination annotation, with Fleiss’ Kappa = 0.857. This serves as a sanity check that the judges are broadly consistent on that binary label and supports the stability of the offline evaluation pipeline. We emphasize, however, that the prompt-ranking score used in Exp. 6 is an aggregate judge score that also depends on additional structured judge fields (see Supplementary Material), so Fleiss’ Kappa should not be interpreted as a direct reliability coefficient for that full aggregate score.
Overall, the results of Exp. 6 suggest that Risk-Cos can serve as a useful, decoding-free criterion for selecting among semantically equivalent prompt formulations. More importantly, this finding provides direct empirical support for the NUP account developed in this paper: within the prefill stage, stronger prompt–gradient coupling is associated with better-constrained continuations, whereas weaker coupling reflects the under-conditioning regime predicted by Proposition 4.
@@ VI-B Unified interpretation across vision and language
Within this limit-based view, adversarial fragility and hallucination become two opposite ways of mismanaging the same uncertainty budget. Table IV summarizes how NUP manifests differently across modalities.
TABLE IV: Two regimes from the NUP. The NUP constrains an effective conjugate volume $V=\Delta x\,\Delta p\,\sqrt{1-\rho^{2}}\geq\kappa$, where $\rho$ is the operator correlation between $\hat{x}_{u}$ and $\hat{p}_{u}$ in the RS inequality. Empirically, we diagnose the input–gradient coupling channel entering $\rho$ using the $C$-Probe: in vision we use $|\cos(x,p)|$ (inputs are normalized), while in LLM prompting we use the centered $|\cos(\bar{x},\bar{p})|$.
@@ VI-D Limitations and future directions
Our analysis is most informative when there exists a nontrivial boundary-relevant subset that the loss-phase construction can emphasize. In the language setting, the usefulness of CC-Probe depends on a nontrivial match between benchmark difficulty and model capability: if the benchmark is too easy hallucinations are too rare, if it is too hard failures become nearly ubiquitous, and the probe is therefore most informative on prompts near the model’s effective uncertainty boundary.
In near “pure fitting” regimes—e.g., when the loss is essentially small on the relevant support, or the input–output mapping behaves close to bijective with little effective compression [40]—the weighting $w_{c}(x)\propto\mathcal{L}_{c}(x)^{2}$ provides little contrast and distinct loss-induced states $\psi_{c}$ are not meaningfully separated; consequently, the notion of a stressed boundary layer becomes ill-posed and the CC-Probe inclines to lose discriminative power. In addition, the probe is loss-dependent (CE in vision; shifted prompt NLL in language), and Exp. 3 indicates robustness can differ across attack losses (CE vs. DLR), motivating multi-loss evaluation (e.g., APGD-CE and APGD-DLR) and defenses that cover multiple geometry channels. Finally, CC-Probe is a proxy rather than a direct estimator of $\rho_{c}(u)$ or slack $S_{c}(u)$, and prefill-only LLM use still requires one backward pass on prompt embeddings. Future work includes designing alternative state constructions when $\mathcal{L}_{c}$ is near-degenerate, developing stable estimators of direction-dependent slack, and extending coupling-based monitoring/control to multimodal and preference-optimized settings.
Crucially, the NUP framework also serves as a predictive warning for future LLM alignment. Currently, LLM hallucinations are predominantly driven by a high-slack (under-conditioned) regime, as detected by the CC-Probe. However, if future training paradigms aggressively reduce this slack to suppress hallucinations—forcing prompts into tight, deterministic couplings with specific continuations—the NUP inequality $\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}\geq\tfrac{1}{2}$ dictates that the system could inevitably shift toward the opposite extreme. In doing so, we risk inducing the same boundary-stress regime currently plaguing vision models, which would manifest in LLMs as extreme adversarial fragility (e.g., imperceptible prompt perturbations or synonymous substitutions triggering catastrophic output shifts). Ultimately, the NUP demonstrates that we cannot entirely eliminate boundary uncertainty; we can only trade one form of fragility for another. Sustainable model design might therefore navigate this fundamental trade-off, seeking an intermediate “Goldilocks” zone rather than minimizing a single failure axis.
@@ VII Conclusion
In this study we have introduced the Neural Uncertainty Principle (NUP), a fundamental geometric constraint showing that input projections and their loss gradients are conjugate observables under a loss-induced state. NUP unifies adversarial vulnerability in vision and hallucination in LLMs as opposite extremes on the same conjugate plane—saturation versus slack. Based on this principle, we derive a practical single-backward probing method, CC-Probe, which enables the diagnosis of cross-modal boundary anomalies and supports targeted interventions: ConjMask and LogitReg for enhancing robustness in vision tasks, and prefill-stage risk scoring for hallucination detection and prompt selection in language models. By reframing two isolated reliability challenges as a single conjugate trade-off, the NUP provides a unified perspective for advancing robust and faithful AI.
@@
Acknowledgments
During manuscript preparation, the authors used ChatGPT (OpenAI) for language editing, cross-section consistency checking, and suggestions aimed at improving the clarity and presentation of mathematical exposition. All technical content, including the underlying proofs, derivations, experimental design, implementation, and conclusions, was developed and verified by the authors, who take full responsibility for the final manuscript.